# 4DGen: Grounded 4D Content Generation with Spatial-temporal Consistency

### v1

@@ Abstract
Aided by text-to-image and text-to-video diffusion models, existing 4D content creation pipelines utilize score distillation sampling to optimize the entire dynamic 3D scene. However, as these pipelines generate 4D content from text or image inputs, they incur significant time and effort in prompt engineering through trial and error. This work introduces 4DGen, a novel, holistic framework for grounded 4D content creation that decomposes the 4D generation task into multiple stages. We identify static 3D assets and monocular video sequences as key components in constructing the 4D content. Our pipeline facilitates conditional 4D generation, enabling users to specify geometry (3D assets) and motion (monocular videos), thus offering superior control over content creation. Furthermore, we construct our 4D representation using dynamic 3D Gaussians, which permits efficient, high-resolution supervision through rendering during training, thereby facilitating high-quality 4D generation. Additionally, we employ spatial-temporal pseudo labels on anchor frames, along with seamless consistency priors implemented through 3D-aware score distillation sampling and smoothness regularizations. Compared to existing baselines, our approach yields competitive results in faithfully reconstructing input signals and realistically inferring renderings from novel viewpoints and timesteps. Most importantly, our method supports grounded generation, offering users enhanced control, a feature difficult to achieve with previous methods. Project page: https://vita-group.github.io/4DGen/
@@ 1 Introduction
Figure 2: Previous work generates 4D content in one click. Our work introduces Grounded 4D Content Generation, which employs a video sequence and an optional 3D asset to specify the appearance and motion.
@@
Despite the exciting progress on either 3D or video generation, little attention has been put to the intersection of these two directions, i.e. the generation of 4D (dynamic 3D) assets, mainly due to the lack of high-quality data. Existing approaches mainly focus on category-specific generation [78, 19, 26, 86, 24].The most related work for general object generation is MAV3D [62], where SDS is applied on text-to-image and text-to-video diffusion models to distill a 4D representation from scratch. However, their results suffer from multiple issues. Firstly, the results are produced from a vague single image or text input, and there is no guarantee that the time-consuming optimization process will lead to the desired 4D result. This introduces extra trial and error costs on prompt engineering in practical applications. Another issue arises from the 4D representation. Though their Hexplane [7] is more efficient than vanilla NeRF, rendering high-resolution images across long frames is still memory-costly and not affordable. As a result, score distillation is conducted at a lower resolution, introducing unpleasant artifacts. Though MAV3D [62] incorporates additional super-resolution models to improve the appearance, the generation is still blurry in texture and has flickering geometry.
To overcome the above issues, we introduce 4DGen, a novel pipeline tackling a new task of Grounded 4D Generation. As shown in Fig. 2, we provide users with explicit fine-grained controllability over the generation of 4D content. Our key idea is to leverage a monocular video and an optional static 3D asset as conditional signals to specify the motion and appearance of the 4D content. By conditioning on the video sequence, explicit motion information is introduced, enabling the user to directly adjust the 4D content in terms of motion. An optional static 3D asset can be provided to adjust the geometry and appearance of the final result, further improving the user experience. When the 3D asset is not specified, our pipeline performs single-image-to-3D generation as a surrogate. When the monocular video sequence is not provided, we leverage existing video generation tools to produce a video sequence.
Our framework introduces novel designs to accomplish the grounded 4D generation task. We implement our 4D asset with deformable 3D Gaussians, an efficient 4D representation that enables high-resolution rendering at the training stage. We first construct static 3D Gaussians to support deformation into motion sequences effectively. Then, we utilize spatial-temporal pseudo labels defined in 2D and 3D space to inject motion information into the 4D representation through anchor frames. Further, we leverage seamless spatial-temporal consistency priors to refine the renderings from arbitrary viewpoints and arbitrary timesteps.
@@
  • We present 4DGen, a holistic pipeline for grounded 4D content generation. Our framework allows full control over the 3D asset’s appearance and motion by specifying a static 3D asset or a monocular video sequence.
  • With the help of the efficient dynamic 3D Gaussian Splatting as scene representation, our model is trained at a high resolution and long frame length, leading to visually pleasing 4D generation, as shown in Fig. 1.
  • Armed with spatial-temporal pseudo labels defined in 2D and 3D space, we explicitly inject motion and appearance information at anchor frames. Moreover, we adopt seamless spatial-temporal consistency priors for arbitrary viewpoint and timestep with the help of 3D-aware score distillation sampling and unsupervised smoothness regularization.
  • Experiments demonstrate our superiority against per-frame generation baselines. Our 4DGen framework delivers faithful reconstruction of the input signals while synthesizing plausible results for novel viewpoints and timesteps.
@@ 2 Related Works
Generating 3D content from multi-modal input has attracted great research interest for years. The task of text-to-3D generation focuses on generating a 3D model from a textual prompt. Due to the lack of large-scale 3D data, many works try to improve 3D generation quality by improving the rendering quality from random 2D viewpoints. Early works DreamFields [25] and CLIPMesh [44] utilized CLIP [51], but the results tend to lack realism. With the help of 2D diffusion priors, the seminal work DreamFusion [49] introduces Score Distillation Sampling (SDS) and showcases encouraging results. Many follow-up works have then tried to improve SDS sampling results through better optimization [31, 23, 41, 21, 4, 72], crafting different 3D representations [34, 63, 81, 10] or constructing better diffusion priors that are suitable for 3D generation [37, 50, 57, 32]. SDS can also be adapted to image-to-3D generation [77, 42, 12, 37]. More recently, with the help of large-scale 3D asset datasets [11, 76, 82], better generative models [37, 38, 27, 59, 36, 22, 79] are built and produce outstanding 3D asset generation ability when given a single image input. Our work draws inspiration from both directions and proposes to use pseudo labels explicitly sampled from a large generative model to supervise anchor viewpoints and employ score distillation sampling to enforce consistency for arbitrary viewpoints.
@@
2.3 Dynamic Content Generation
Unlike static content generation, producing high-quality dynamic datasets remains challenging due to the lack of data [71, 5]. Many prior works seek to leverage existing priors learned on large-scale text-image pairs [56, 6, 9] and adapt to dynamic content generations. Tune-a-Video [75] tackles one-shot video generation through subtle architecture modifications and sub-network tuning. Text2Video Zero [30] introduces a training-free method for animating a pre-trained text-to-image model. AnimateDiff [18] converts existing text-to-image diffusion models to video generators by designing a motion modeling module that is zero-shot transferrable to unseen stable diffusion variants.
Regarding 4D content generation, early works mainly focus on category-specific generation, such as digital humans [78, 19, 24, 39, 33, 67, 84, 53], animals [26] and flowers [86]. Regarding general object generation, MAV3D [62] leverages score distillation sampling from both image and video diffusion models to optimize a 4D scene. Consistent4D [28] is a concurrent method that also studies lifting a monocular video sequence to 4D. Their work leverages frame interpolation-driven consistency loss and per-frame SDS from a 3D-aware diffusion model. Our work similarly studies utilizing 3D-aware diffusion priors for 4D generation, but differs in that we not only employ SDS supervision but also introduce DDIM-sampled pseudo labels for supervision at a higher resolution. Moreover, while their work adopts costly NeRF-based volumetric rendering, we utilize an efficient rasterizer brought by 3D Gaussians, supporting faster training implemented through rendering losses.
3 4DGen: Grounded 4D Generation
Generating 4D contents without 4D annotation is highly ill-posed since multiple 4D results can all be plausible when projecting to a monocular video sequence or a static 3D asset. To this end, we adopt a pre-trained 3D-aware multi-view diffusion model to provide spatial-temporal pseudo labels. Additionally, since the spatial-temporal pseudo labels are only provided on limited 2D renderings, we include score distillation sampling from a 3D-aware diffusion model to supervise random viewpoints and unsupervised smoothness regularization to enforce temporal smoothness at random timestep in 3D explicitly.
3.1 Background
Score Distillation Sampling
Score Distillation Sampling (SDS) is widely adopted for distilling 2D image priors from a pre-trained diffusion model $\epsilon_{\phi}$ into 3D representations through optimization on the renderings. When provided with a 3D model parameterized by $\theta$ and an image generator $g$, the multi-view image $x$ can be rendered by $g(\theta)$.
SDS optimizes the 3D parameter $\theta$ as follows:
@@ 3.2 Static Gaussian Generation
When the user specifies a 3D asset, we can readily construct the static scene by converting the specified 3D asset into a point cloud and subsequently initializing 3D Gaussians from this point cloud. However, in the absence of a static 3D asset, our framework must generate the static Gaussian initialization to enable scene deformation. Owing to the ill-posed and data-limited nature of our generative task, additional design considerations are necessary to construct a plausible static scene.
To achieve this, we utilize a multi-view diffusion model, pre-trained on 3D datasets, to infer the scene structure. We perform DDIM sampling using SyncDreamer [38] to obtain multi-view predictions of the scene, and then optimize a set of static 3D Gaussians using the sampled images. Our framework facilitates both random initialization and initialization from pre-processed point clouds for the 3D Gaussians. Specifically, for the initial frame, off-the-shelf multi-view reconstruction algorithms [68, 70] are employed to estimate coarse geometry and extract point clouds.
@@ 3.3 Spatial-temporal Anchor Frame Pseudo-labels
Spatial-temporal 2D Pseudo Labels
@@
3D Deformation Pseudo Label
Since these supervisions are defined on the 2D renderings, we rely on the differentiable rendering to propagate essential knowledge to the 3D points. However, it is observed that when Gaussians are presented in incorrect locations, they tend to reduce their visibility by changing their appearance instead of moving to the desired locations. Though ideally, this issue can be alleviated through scene flow supervision, estimating these dense 3D point correspondences from monocular video is highly challenging and requires a time-consuming optimization process [69].
To this end, we propose constructing point clouds from SyncDreamer predictions and treating them as pseudo labels for scene deformation. Note that point clouds at different frames are unrelated to 3D point correspondences because the points constructing the point clouds are stored as unordered sets and constructed individually. Therefore, we use Chamfer distance for the 3D Gaussian locations, which measures the difference between two point sets without the need of point correspondences. It instead searches for the nearest neighbor as an alternate, defined as follows,
$\mathcal{L}_{\text{chamfer}}$ $=\frac{w_{1}}{\left|P_{1}\right|}\sum_{p_{1i}\in P_{1}}\min_{p_{2j}\in P_{2}}\left(\left\|p_{1i}-p_{2j}\right\|_{2}^{2}\right)$ (4)
$+\frac{w_{2}}{\left|P_{2}\right|}\sum_{p_{2j}\in P_{2}}\min_{p_{1i}\in P_{1}}\left(\left\|p_{2j}-p_{1i}\right\|_{2}^{2}\right),$
where $P_{1}$ is our 3D Gaussians 3D mean and $P_{2}$ refers to the pre-processed point clouds at each timeframe. This novel design effectively prevents the scene from mode-collapsing into static, allowing us to integrate the framework with score distillation sampling [49], which is widely observed to be prone to mode collapse. The loss functions for pseudo labels are summarized as follows,
$\mathcal{L}_{\text{pseudo}}=\mathcal{L}_{\text{sync}}+\omega_{1}\mathcal{L}_{\text{chamfer}},$ (5)
where $\omega_{1}$ is a weighting factor and $\mathcal{L}_{\text{sync}}$ refers to the LPIPS [85] distance between the rendered output and the pseudo label.
@@ 3.4 Seamless Spatial-temporal Consistency Priors
Arbitrary Viewpoint Spatial Consistency through SDS
@@
Arbitrary Timestep Temporal Consistency via Smoothness Regularization
@@
$\mathcal{L}_{\text{TV}}(\mathbf{P})$ $=\frac{1}{|C|n^{2}}\sum_{c,i,j}\left(\|\mathbf{P}_{c}^{i,j}-\mathbf{P}_{c}^{i-1,j}\right\|_{2}^{2}$ (6)
$+\left\|\mathbf{P}_{c}^{i,j}-\mathbf{P}_{c}^{i,j-1}\right\|_{2}^{2}),$
@@
Figure 4: Comparison with baseline methods. We provide video comparisons in supplementary. The reference only contains front-view images, while our 4DGen model is capable of rendering at any timestep from any viewpoint. Our method maintains greater consistency and higher quality in both spatial and temporal consistency. In contrast, other methods result in significant artifacts or blurring, attributed to inaccurate feature estimation from different views. Moreover, these baselines generate individual frames separately and cannot be rendered at intermediate timesteps.
@@ 4 Experiments
Figure 5: More cases about results of 4DGen. Our method achieves high-quality detail from arbitrary viewpoints and timesteps, meanwhile maintains seamless spatial-temporal consistency.
@@ 4.1 Implementation Details
All experiments are conducted using a single RTX 3090 GPU. The training time is highly effected by the number of 3D Gaussians. The time is about 45 minutes with around 25,000 3D Gaussians and 120 minutes with about 90,000 3D Gaussians. Densification introduces more 3D Gaussians, increasing the training time and also bringing better details.
@@ 4.2 Baseline Methods
MAV3D [62] is the sole work preceding 4D generation for 4D content creation. A direct comparison is not feasible since their code has not been released. To evaluate our method, we instead establish two baseline approaches following the practice in MAV3D. DreamGaussian [63] is a state-of-the-art image-to-3D method utilizing 3D Gaussian splatting. DreamGaussian is provided with individual video frames to generate a total of $T$ 3D models, where $T$ is the total number of frames. Subsequently, we can render images from every view at each time. For a fair comparison, we only use their first stage, which optimizes 3D Gaussians directly. The second baseline involves One-2-3-45 [36], a high-quality multi-view diffusion model. Given an image at each time, One-2-3-45 can generate multi-view images and fuse them into a mesh. For these two baselines, we concatenate the $T$ results to construct a 4D model. We also provide comparisons against a concurrent method Consistent4D [28]. The datasets consist of three parts, with each test case containing only a reference video sequence from the front view. The first part consists of a video dataset released by Consistent4D [28]. The second part of the dataset is collected either in the wild or from Sketchfab [1] by ourselves. The final part is generated using a pre-trained image-to-video diffusion model  [3].

### v3

@@ Abstract
Aided by text-to-image and text-to-video diffusion models, existing 4D content creation pipelines utilize score distillation sampling to optimize the entire dynamic 3D scene. However, as these pipelines generate 4D content from text or image inputs directly, they are constrained by limited motion capabilities and depend on unreliable prompt engineering for desired results. To address these problems, this work introduces 4DGen, a novel framework for grounded 4D content creation. We identify monocular video sequences as a key component in constructing the 4D content. Our pipeline facilitates controllable 4D generation, enabling users to specify the motion via monocular video or adopt image-to-video generations, thus offering superior control over content creation. Furthermore, we construct our 4D representation using dynamic 3D Gaussians, which permits efficient, high-resolution supervision through rendering during training, thereby facilitating high-quality 4D generation. Additionally, we employ spatial-temporal pseudo labels on anchor frames, along with seamless consistency priors implemented through 3D-aware score distillation sampling and smoothness regularizations. Compared to existing video-to-4D baselines, our approach yields superior results in faithfully reconstructing input signals and realistically inferring renderings from novel viewpoints and timesteps. More importantly, compared to previous image-to-4D and text-to-4D works, 4DGen supports grounded generation, offering users enhanced control and improved motion generation capabilities, a feature difficult to achieve with previous methods. Project page: https://vita-group.github.io/4DGen/
@@ 1 Introduction
Despite the exciting progress on either 3D or video generation, little attention has been put to the intersection of these two directions, i.e. the generation of 4D (dynamic 3D) assets, like Fig. 1, mainly due to the lack of high-quality data. Existing approaches mainly focus on category-specific generation [83, 22, 31, 93, 28]. The most related work for general object generation is MAV3D [68], where SDS is applied on text-to-image and text-to-video diffusion models to distill a 4D representation from scratch. Consistent4D [32] uses interpolation-driven consistency loss and per-frame SDS to lift a monocular video sequence to 4D. However, their results suffer from multiple issues. Firstly, MAV3D’s reliance on ambiguous single image or textual inputs does not reliably yield desired 4D outcomes with coherent motion, often resulting in limited motion scope and necessitating extensive prompt engineering, thus increasing trial and error costs in practical scenarios. Secondly, the Hexplane [9] representation, despite its efficiency over traditional NeRF approaches, struggles with memory-intensive, high-resolution rendering over extended frames, compelling the use of lower resolution in score distillation and leading to artifacts. Though MAV3D [68] and Consistent4D [32] attempt to mitigate these issues through super-resolution models, the outcomes frequently exhibit blurred textures and inconsistent geometry, detracting from the overall quality of the generated 4D assets.
To address the aforementioned challenges, we introduce 4DGen, a novel pipeline tackling a new task of Grounded 4D Generation, which focuses on video-to-4D generation. As shown in Fig. 2, our primary strategy involves using monocular videos as conditional inputs to provide users with precise control over both the motion and appearance of generated 4D content. This strategy addresses the limited motion capabilities and controllability issues associated with previous methods. Notably, videos can be either user-provided or synthesized through video diffusion processes. With the support of an image-to-video diffusion model [3], our pipeline facilitates an image-to-video-to-4D conversion.
Additionally, we implement our 4D asset with deformable 3D Gaussians, an efficient 4D representation that enables high-resolution rendering at the training stage. Specifically, we first construct static 3D Gaussians to support deformation into motion sequences effectively. Subsequently, we employ spatial-temporal pseudo labels generated by a multi-view diffusion model and score distillation sampling loss. This approach injects motion information into the 4D representation at anchor frames. Furthermore, we use seamless spatial-temporal consistency priors to refine renderings from any viewpoint and at any timestep.
@@
  • We present 4DGen, a novel pipeline for grounded 4D content generation. Our framework allows full control over the 4D asset’s appearance and motion by specifying a monocular video sequence.
  • Leveraging efficient dynamic 3D Gaussian Splatting for scene representation, our model is trained at a high resolution and long frame length, leading to visually pleasing 4D generation, as shown in Fig. 1.
  • Utilizing spatial-temporal pseudo labels, we directly incorporate motion and appearance information into anchor frames. Additionally, we ensure seamless spatial-temporal consistency from any viewpoint and at any timestep through 3D-aware score distillation sampling and unsupervised smoothness regularization.
Our experiments showcase 4DGen’s superiority over per-frame generation baselines and existing 4D generation techniques in video-to-4D tasks. Our 4DGen framework delivers faithful generation of the input signals while synthesizing plausible results for novel viewpoints and timesteps. Compared with previous image-to-4D and text-to-4D works, 4DGen offers better motion generation and increased user control, surpassing the capabilities of prior methods.
Figure 2: Previous work generates 4D content by vague single image or text input, leading to uncertain results. Our work introduces Grounded 4D Content Generation, which employs a video sequence specify the appearance and motion. Our work supports video-to-4D generation, where the input video can either be user-specified or generated from video diffusion models.
@@ 2 Related Works
Diffusion models are widely used across various fields [60, 54, 92, 87, 26, 27], and the generation of 3D content from multi-modal inputs has garnered significant research interest over the years. The task of text-to-3D generation focuses on generating a 3D model from a textual prompt. Due to the lack of large-scale 3D data, many works try to improve 3D generation quality by improving the rendering quality from random 2D viewpoints. Early works DreamFields [30] and CLIPMesh [49] utilized CLIP [57], but the results tend to lack realism. With the help of 2D diffusion priors, the seminal work DreamFusion [55] introduces Score Distillation Sampling (SDS) and showcases encouraging results. Many follow-up works have then tried to improve SDS sampling results through better optimization [36, 29, 46, 24, 5, 77, 95], crafting different 3D representations [39, 69, 86, 12] or constructing better diffusion priors that are suitable for 3D generation [42, 56, 63, 37]. SDS can also be adapted to image-to-3D generation [82, 47, 14, 42]. More recently, with the help of large-scale 3D asset datasets [13, 81, 88], better generative models [42, 43, 33, 65, 41, 25, 84] are built and produce outstanding 3D asset generation ability when given a single image input. Our work draws inspiration from both directions and proposes to use pseudo labels explicitly sampled from a large generative model to supervise anchor viewpoints and employ score distillation sampling to enforce consistency for arbitrary viewpoints.
2.3 Dynamic Content Generation
Unlike static content generation, producing high-quality dynamic datasets remains challenging due to the lack of data [75, 7]. Many prior works seek to leverage existing priors learned on large-scale text-image pairs [62, 8, 11] and adapt to dynamic content generations. Tune-a-Video [80] tackles one-shot video generation through subtle architecture modifications and sub-network tuning. Text2Video-Zero [35] introduces a training-free method for animating a pre-trained text-to-image model. AnimateDiff [21] converts existing text-to-image diffusion models to video generators by designing a motion modeling module that is zero-shot transferrable to unseen stable diffusion variants.
Regarding 4D content generation, early works mainly focus on category-specific generation, such as digital humans [83, 22, 28, 44, 38, 73, 90, 59], animals [31] and flowers [93]. Regarding general object generation, MAV3D [68] leverages score distillation sampling from both image and video diffusion models to optimize a 4D scene. Consistent4D [32] is a concurrent method that also studies lifting a monocular video sequence to 4D. Their work leverages frame interpolation-driven consistency loss and per-frame SDS from a 3D-aware diffusion model. Animate124 [94] and 4D-fy [6] are concurrent methods achieve image-to-4D or text-to-4D by video SDS loss. Our work similarly studies utilizing 3D-aware diffusion priors for 4D generation, but differs in that we not only employ SDS supervision but also introduce DDIM-sampled pseudo labels for supervision at a higher resolution. Moreover, while their work adopts costly NeRF-based volumetric rendering, we utilize an efficient rasterizer brought by 3D Gaussians, supporting faster training implemented through rendering losses.
3 4DGen: Grounded 4D Generation
Generating 4D contents without 4D annotation is highly ill-posed since multiple 4D results can all be plausible when projecting to a monocular video sequence. To this end, we adopt a pre-trained 3D-aware multi-view diffusion model to provide spatial-temporal pseudo labels. Additionally, since the spatial-temporal pseudo labels are only provided on limited 2D renderings, we include score distillation sampling from a 3D-aware diffusion model to supervise random viewpoints and unsupervised smoothness regularization to enforce temporal smoothness at random timestep in 3D explicitly.
@@ 2.3 Dynamic Content Generation
3.1 Preliminary
3.1.1 Score Distillation Sampling
Score Distillation Sampling (SDS) is widely adopted for distilling 2D image priors from a pre-trained diffusion model $\epsilon_{\phi}$ into 3D representations through optimization on the renderings. When provided with a 3D model parameterized by $\theta$ and an image generator $g$, the multi-view image $x$ can be rendered by $g(\theta)$. SDS optimizes the 3D parameter $\theta$ as follows:
@@ 3.2 Static Gaussian Generation
Our framework initially generates the static Gaussian initialization to enable scene deformation. Owing to the ill-posed and data-limited nature of our generative task, additional design considerations are necessary to construct a plausible static scene. To achieve this, we utilize a multi-view diffusion model, pre-trained on 3D datasets, to infer the scene structure. We perform DDIM sampling using SyncDreamer [43] to obtain multi-view predictions of the scene, and then optimize a set of static 3D Gaussians using the sampled images. Our framework adopts random initialization and optimizes via $\mathcal{L}_{1}$ loss, unlike using SDS loss in DreamGaussian [69]. Specifically, for the initial frame, off-the-shelf multi-view reconstruction algorithms [74, 76] are employed to estimate coarse geometry and extract point clouds.
@@ 3.3 Spatial-temporal Anchor Frame Pseudo-labels
3.3.1 Spatial-temporal 2D Pseudo Labels
@@ 3.4 Seamless Spatial-temporal Consistency Priors
3.4.1 Arbitrary Viewpoint Spatial Consistency through SDS
@@
3.4.2 Arbitrary Timestep Temporal Consistency via Smoothness Regularization
@@
$\mathcal{L}_{\text{TV}}(\mathbf{P})$ $=\frac{1}{|C|n^{2}}\sum_{c,i,j}\left(\|\mathbf{P}_{c}^{i,j}-\mathbf{P}_{c}^{i-1,j}\right\|_{2}^{2}+\left\|\mathbf{P}_{c}^{i,j}-\mathbf{P}_{c}^{i,j-1}\right\|_{2}^{2}),$ (4)
@@ 4.2 Baseline Methods
MAV3D [68] is the first work preceding 4D generation for 4D content creation. A direct comparison is not feasible since their code has not been released. To evaluate our method, we instead establish two baseline approaches following the practice in MAV3D. DreamGaussian [69] is a state-of-the-art image-to-3D method utilizing 3D Gaussian splatting. DreamGaussian is provided with individual video frames to generate a total of $T$ 3D models, where $T$ is the total number of frames. Subsequently, we can render images from every view at each time. The second baseline involves One-2-3-45 [41], a high-quality multi-view diffusion model. Given an image at each time, One-2-3-45 can generate multi-view images and fuse them into a mesh. For these two baselines, we concatenate the $T$ results to construct a 4D model.
We also provide comparisons against a concurrent video-to-4D method Consistent4D [32], image-to-4D method Animate124 [94] and text-to-4D method 4D-fy [6]. The datasets consist of three parts, with each test case containing only a reference video sequence from the front view. The first part consists of a video dataset released by Consistent4D [32]. The second part of the dataset is collected either in the wild or from Sketchfab [1] by ourselves. The final part is generated using a pre-trained image-to-video diffusion model  [3].
Figure 4: Comparison with baseline methods. The reference only contains front-view images, while our 4DGen model is capable of rendering at any timestep from any viewpoint. Our method maintains greater consistency and higher quality in both spatial and temporal consistency. In contrast, other methods result in significant artifacts or blurring, attributed to inaccurate feature estimation from different views. Moreover, these baselines generate individual frames separately and cannot be rendered at intermediate timesteps. We provide more video comparisons in the project page.
Figure 5: Image-to-4D Comparison. Our input video sequences are generated through the process image-to-video via Stable Video Diffusion [3]. As shown in bounding box areas, our results exhibit identity consistency with the input reference image when compared to Animate124. We provide video comparisons in the project page.